### v2

# A Survey of Large Language Model Empowered Agents for Recommendation and Search: Towards Next-Generation Information Retrieval

Abstract.
Information technology has profoundly altered the way humans interact with information. The vast amount of content created, shared, and disseminated online has made it increasingly difficult to access relevant information. Over the past two decades, recommender systems and search (collectively referred to as information retrieval systems) have evolved significantly to address these challenges. Recent advances in large language models (LLMs) have demonstrated capabilities that surpass human performance in various language-related tasks and exhibit general understanding, reasoning, and decision-making abilities. This paper explores the transformative potential of LLM agents in enhancing recommender and search systems. We discuss the motivations and roles of LLM agents, and establish a classification framework to elaborate on the existing research. We highlight the immense potential of LLM agents in addressing current challenges in recommendation and search, providing insights into future research directions. This paper is the first to systematically review and classify the research on LLM agents in these domains, offering a novel perspective on leveraging this advanced AI technology for information retrieval. To help understand the existing works, we list the existing papers on LLM agent based recommendation and search at this link: https://github.com/tsinghua-fib-lab/LLM-Agent-for-Recommendation-and-Search.
Large language model agent; recommender system; search system; information system
journal: TOIS
1. Introduction
Information technology has fundamentally transformed human life, particularly in how we interact with information. The development of the information age has led to the creation, sharing, and dissemination of vast amounts of content on the internet, making it increasingly challenging for individuals to access relevant information. Against this backdrop, the past two decades have witnessed the evolution of information systems, including Recommender Systems (RSs) (Zhang et al., 2019; Gao et al., 2023b) and search systems (Brin and Page, 1998; Speretta and Gauch, 2005). Typically, search systems passively retrieve relevant content from vast amounts of information with users’ query words, while recommender systems actively generate a list of candidates that users might be interested in based on collected user behavioral data or profile information (Schütze et al., 2008; Aggarwal et al., 2016; Levene, 2011). Some researchers consider recommendation as a specialized form of search, where the understanding of potential user needs is modeled as search terms (Xu et al., 2018).
Recently, large language models (LLMs) have achieved remarkable success and are considered one of the most viable paths toward general artificial intelligence (Achiam et al., 2023). LLMs have demonstrated capabilities similar to or even surpass human performance, not just in language-related tasks, exhibiting general understanding, reasoning, and decision-making abilities (Zhao et al., 2023b). Researchers have leveraged the text processing and commonsense reasoning capabilities of large models to conduct a series of recommendation and search tasks (White, 2024; Zhu et al., 2023; Ai et al., 2023). Furthermore, to overcome the limitations of LLMs, researchers build various LLM agents (Xi et al., 2023). Specifically, LLM agents use LLM as the core and are equipped with functions such as memory management, workflow, input-output interfaces, external tool invocation, and large-small model collaboration, making their intelligence levels closer to human and capable of handling more complex tasks (Zhao et al., 2024a; Park et al., 2023; Zhou et al., 2023c).
For the crucial and fundamental tasks of recommendation and search, researchers increasingly recognize the high value of LLM agents. In this paper, we make the first attempt to provide a systematic and comprehensive review of the research efforts on recommendation and search with LLM agents. We first introduce relevant background knowledge and further discuss the motivations for utilizing LLM agents, as well as how LLM agents address those critical challenges faced by recommendation and search. Next, we establish a taxonomy, categorizing these recent advances according to the role of LLM agents in the overall recommender or search system. We then elaborate on these works and analyze how they construct and use LLM agents. It is worth mentioning there have already been several works discussing the application of LLMs in the fields of IR, which either focus solely on LLMs (Chen et al., 2024a; Bao et al., 2024; Wu et al., 2024b; Gao et al., 2024) or a specific IR application (Peng et al., 2025), but this survey paper significantly differs by focusing on the application of LLM agents across the entire spectrum of IR functionalities.
The contributions of this paper can be summarized as follows:
  • First, this paper is the first to review and organize the research on LLM agents in the research of search and recommendation, which is a new yet fast-growing research field.
  • Second, we construct a taxonomy that well organizes the existing work by answering the fundamental question of why LLM agents are needed and how LLM agents enhance recommendation and search.
  • Finally, we conduct open discussions about the unresolved challenges and important future research directions, which can inspire the following works in this area.
The structure of this paper is as follows: In Section 2, we introduce the background knowledge of recommender systems, search systems, LLMs, and LLM agents. In Section 3, we analyze why LLM agents are necessary for recommendation and search. In Section 4, we first establish a coherent and comprehensive taxonomy system and then elaborate on the existing work. In Section 5, we extensively discuss the advancements of embodied LLM agents within the cyber environment, highlighting their potential to inspire the next generation of recommendation and search. In Section 6, we systematically analyze the current problems and significant future directions for recommendation and search based on LLM agents. Finally, we conclude the paper in Section 7.
2. Background
2.1. Recommendation and Search
Recommendation and search are pivotal challenges in information retrieval and have been extensively studied for decades. Despite their distinct application scenarios, they share similar core components (namely Interaction Interface, User/Query Modeling, Item Modeling, and Matching, Ranking, and Re-ranking). This section introduces the background and concepts of each core component.
2.1.1. Interaction Interface
The rapid development of smart devices and the mobile Internet has exposed users to numerous applications employing recommender systems and search engines. Interacting with a traditional recommender system is simple, where users are mostly unconsciously browsing contents (e.g., on an e-commerce website, or a video sharing platform) provided by the recommender system (Resnick and Varian, 1997). On the other hand, the interaction interface between users and search engines usually consists of a keyword query submission from users (e.g., on a search engine website or a project management software) and a one-round result retrieval from the search engines (Baeza-Yates et al., 1999). Conversational recommender systems and search engines have recently gained attention (Zhang et al., 2018). In this interface, users interact through a chat interface using natural language, iteratively providing additional information or feedback based on previous results, forming a multi-turn dialogue. This conversational interface is more closely aligned with human communication methods and achieves more precise and satisfying results.
2.1.2. User/Query Modeling
For recommender systems and search engines to retrieve satisfying results, they must understand the users’ intentions. For example, identifying the products a user wants to buy on an e-commerce app or determining relevant websites based on a keyword query. This requires a precise modeling of user characteristics and input queries. In recommender systems, this includes utilizing the user profile (gender, age, preferences, etc.), browsing/clicking/purchasing history, and more (Lops et al., 2011). In search engines, this involves text segmentation, named entity recognition, part-of-speech tagging, relation extraction, etc (Croft et al., 2010). Advanced systems like multi-modal and conversational search engines also incorporate image recognition, multi-turn dialogue comprehension, and other algorithms (Jannach et al., 2021; Wang et al., 2016).
2.1.3. Item Modeling
Apart from user/query modeling, item modeling also plays an essential role in efficient and extensive retrieval. It gathers information of the items, and preprocesses them into some representation (e.g., a feature vector) for later usage. In recommender systems, item modeling considers item descriptions, attributes, images, buyers, and other related information (Lops et al., 2011; Liu et al., 2024). In traditional search engines, this process is known as Indexing (Croft et al., 2010), where significant keywords or tags from the content of web pages are identified and stored beforehand for faster retrieval. Advanced search engines may also use auxiliary knowledge bases specific to certain domains.
2.1.4. Matching, Ranking, and Re-ranking
To effectively retrieve accurate items according to the users/queries from a huge-size candidate item set (usually millions or more), a recommender system or a search engine typically comprises of three consecutive stages: Matching, Ranking, and Re-ranking (Hron et al., 2021).
  • Matching: The Matching (or Recall) (Covington et al., 2016) stage retrieves potentially relevant hundreds of items from the full millions of candidates. In this stage, the algorithm is usually coarse (e.g., vector inner product) and highly efficient due to the huge-size input item set.
  • Ranking: The Ranking (Cheng et al., 2016) stage ranks the output of the Matching stage and selects the top ten of the items. With a smaller item set, this stage employs more complex models (e.g., deep neural networks) for higher accuracy.
  • Re-ranking: The Re-ranking (Pei et al., 2019) stage further adjusts the final item list according to some additional requirements such as diversity, fairness, freshness, and business goals, etc.
By leveraging these stages, recommender systems and search engines ensure precise and relevant results.
2.2. Large Language Model and LLM Agents
2.2.1. Large Language Models
Language models play a pivotal role in Natural Language Processing (NLP), primarily used for assessing sentence probabilities and generating grammatically correct text. In its early stages, the development of language models relied on statistical and probability theories such as n-grams, maximum entropy, and Hidden Markov Models, but was limited by its ability to handle long-range dependencies, known as the “curse of dimensionality.” With the rise of deep learning technologies, neural network language models gradually became mainstream due to their robust generalization capabilities and flexible architectural designs. Models like Recurrent Neural Networks (RNNs) (Sherstinsky, 2020), Long Short-Term Memory networks (LSTMs) (Graves and Graves, 2012), and Gated Recurrent Units (GRUs) (Cho et al., 2014) effectively mitigated the challenges of processing long sequences by utilizing internal memory units, thereby better capturing contextual dependencies within sequences and significantly enhancing language model performance.
The introduction of the Transformer model (Vaswani et al., 2017) revolutionized the NLP field by replacing recurrent structures with self-attention mechanisms, thereby improving training speed and parallelism. Concurrently inspired by the pre-training and fine-tuning paradigm from computer vision, language models began leveraging large amounts of unlabeled data for pre-training, followed by task-specific fine-tuning, which gave rise to high-performance models such as BERT (Devlin et al., 2018). With the rapid expansion of training data and significant growth in computational resources, LLMs have established absolute dominance in the NLP. This wave not only symbolizes the comprehensive transformation of NLP from focusing on single-task optimization to multimodal perception and generation but also heralds language models reaching new heights in generation, understanding, and transformation. In the latest research, LLMs such as GPT-4 (Achiam et al., 2023), Claude (Anthropic, 2024), etc., not only perform well in tasks across various NLP tasks such as text generation, multilingual translation, and sentiment analysis but also demonstrate impressive performance in tasks traditionally considered to require deep cognitive abilities like solving mathematical problems, logical reasoning, and strategic planning. This comprehensive performance marks the maturity of artificial intelligence technology, steadily advancing in a direction closer to Artificial General Intelligence (AGI).
2.2.2. LLM Agents
The concept of agent originated in the 1960s, aiming to develop software entities that can simulate human intelligence. By the 1970s and 1980s, the rise of expert systems (Waterman, 1985) represented the ability of intelligent programs to solve professional problems based on fixed rules and knowledge bases. Although these systems were powerful, they were limited by static knowledge and rigid logic. In the late 1980s, researchers such as Michael Wooldridge redefined agents, emphasizing their autonomy, responsiveness to external events, and interaction with the environment and other agents. Minsky elaborated on the concept of agents in his work (Minsky, 1988), describing them as individuals that can participate in social interactions and exhibit intelligent behavior. Minsky’s idea that many relatively simple components (i.e., agents) in a complex system can solve complex problems by collaborating with each other later became the basis for research in Multi-Agent Systems (MAS) (Dorri et al., 2018) and distributed artificial intelligence (Bond and Gasser, 2014). In the 1990s, with the popularity of the Internet and distributed computing, the research focus turned to the cooperation of agents in an open, dynamic environment. This required them to have intelligent strategies and interaction capabilities, promoting complex intelligent ecosystems’ design.
At the beginning of the 21st century, breakthroughs in machine learning, particularly represented by the rise of deep learning, enabled autonomous learning from massive amounts of data rather than relying solely on pre-set programming rules. This marked a shift in agent research from rule-based methods to data-driven approaches. Deep neural networks significantly enhance the capabilities of intelligent agents in areas such as image recognition, speech processing, and natural language understanding, thus ushering in a new era of Artificial Intelligence (AI) agents. Recently, GPT series (OpenAI, 2022; Achiam et al., 2023) and other major language models have made significant progress in NLP, greatly improving the performance of artificial intelligence agents and demonstrating their profound potential.
Figure 1. Schematic diagram of the three core modules of agent
An agent is an autonomous computing entity capable of perceiving its environment, making decisions, and taking actions. Figure 1 shows the three core modules of the LLM agent, each detailed as follows:
  • Perception Module: The perception module is equivalent to the agent’s “senses”, ensuring that the agent can obtain the information needed for decision-making on time. Unlike conventional single-channel information processing systems, LLM agents demonstrate superior capability in seamlessly integrating multimodal information sources, including textual, auditory, visual, and even video data, thereby establishing a more robust and comprehensive perceptual framework (Schumann et al., 2024).
  • Control Module: The control module plays the role of a proxy “brain”, analyzing the collected information, evaluating possible action plans, and selecting the best strategy. In LLM agents, an LLM serves as the core of the control module, enhancing the agent’s understanding and generation ability of language contexts and strengthening its situational adaptability, creativity, and long-term memory functions.
  • Action Module: The action module acts as the agent’s “limbs”, executing decisions to alter environmental states or achieve goals. In the LLM agent, this module is not limited to simple execution actions. Still, it is endowed with deeper functions, which can realize complex tasks such as language generation, tool use, and even concrete actions, greatly expanding the application scope and influence of the agent.
Currently, LLM agents have emerged in many fields. They can not only handle basic transactional work but also involve complex tasks that require deep understanding and creative thinking. For example, in game development, LLM agents can give non-player characters (NPCs) vivid and natural language capabilities to create a more immersive gaming environment (Park et al., 2023; Hu et al., 2024); in the field of education, LLM agents can provide personalized tutoring for students as classroom assistants and help teachers correct homework (Lagakis and Demetriadis, 2024). In the field of scientific research, LLM agents such as CellAgent can perform complex tasks such as single-cell RNA sequencing data analysis that require the combination of specific domain knowledge (Xiao et al., 2024). LLM agents are constantly expanding the boundaries of human capabilities. As the technology matures, more innovative application scenarios will continue to emerge. LLM agents will drive human civilization in a more intelligent and sustainable direction.
3. Why LLM agents can be used for recommendation and search
As discussed in previous sections, LLM agents introduce new capabilities compared to simpler models and have diverse applications, such as game NPCs, education, RNA design, etc. In this section, we will demonstrate why these new capabilities provide LLM agents with distinct advantages in addressing IR problems.
3.1. LLM Agents Can Do Deep Thinking and Task Decomposition
In an era of information overload, user queries are often complex and diverse, making it challenging for traditional information retrieval systems to grasp the users’ true needs accurately. However, LLM agents, with their advanced language comprehension abilities, can thoroughly analyze user queries, identifying key information, implicit intents, and contextual relationships. Two key capabilities underpin LLM agents’ advances: deep thinking and extended information input. Firstly, Deep thinking enables the agent to perform a more in-depth analysis of a given issue, typically achieved through techniques of Chain-of-Thought (CoT) (Wei et al., 2022; Feng et al., 2024). Secondly, the use of extended information inputs enables the agent to remember more content, including user preferences, context information, and item profiles. It is usually facilitated by technologies such as a long context window that supports millions of token inputs. (Ding et al., 2024; Jin et al., 2024a). Together, these capabilities enable the agent to decompose a complex IR task into multiple sub-tasks for more effective execution (Joko et al., 2024; Aliannejadi et al., 2024). For example, when a user plans a trip, an LLM agent can decompose the process into selecting a destination, planning the itinerary, booking flights and hotels, and calculating the budget, ultimately presenting the user with a complete report (Xie and Zou, 2024), which can be quite challenging for traditional IR systems.
3.2. LLM Agents can Interact with Environment and Integrate Information
In the field of information retrieval, another significant advantage of LLM agents is their ability to interact with environments and integrate results (He et al., 2025; Zhou et al., 2023a). IR often involves multiple data sources and complex retrieval scenarios where traditional retrieval systems may struggle to collect information effectively from different sources. LLM agents with action modules can use various tools, browse web pages, operate mobile apps, and even independently search the internet to find needed information (Schick et al., 2024; Qin et al., 2023). Additionally, LLM agents can go a step further by filtering and summarizing the collected content, presenting users with a clear and concise report instead of delivering raw, unprocessed results like traditional IR systems. For example, when conducting academic literature searches, an LLM agent can interact with multiple academic databases to obtain research outcomes from different fields. It then analyzes and consolidates these results to provide users with a comprehensive literature review. Furthermore, LLM agents can interact with users, continuously adjusting retrieval strategies based on user feedback, thereby enhancing the quality of the retrieval results.
3.3. LLM Agents Can Serve as User Simulators to Generate Feedback to Information System
IR systems rely on real user feedback for improvement, but conducting experiments on actual users can be costly and negatively impact user experience. The various modules within LLM agents enable them to simulate human perception, memory, and actions, making them well-suited for collecting feedback in place of real users. LLM agents can simulate real user responses when using a product and identifying potential improvements (Zhang et al., 2024g). They can also be used to explore the evolution of users’ mindsets and interests, aiding researchers in better understanding user behavior (Park et al., 2023; Takata et al., 2024). LLM agents can simulate users’ preferences and interact with recommender and search systems. These generated interaction data can be used to train more robust IR models (Sekulić et al., 2024b; Zhang et al., 2024f).
In summary, LLM agents bring several critical capabilities to the table that make them highly suitable for recommendation and search tasks. Their ability to think and plan, interface with various tools, and simulate user behavior makes a significant difference in the user experience of IR in the new era.
4. Recent Advances of LLM Agents for Recommendation and Search
Figure 2. Illustration of existing work of LLM agents for recommendation and search.
Recently, significant advancements have been made in the research on combining LLM agents with traditional recommendation and search algorithms. This section provides a comprehensive overview of these developments, categorized into the contributions and innovations made in three primary areas: the overall domain and taxonomy, the role of agents, and recent significant papers. The classification of existing work in LLM agents for Recommendation and Search is shown in Figure 2.
4.1. LLM Agents for Recommendation
4.1.1. Taxonomy Introduction
Deploying LLM agents in recommendation and search domains has given rise to new frameworks and methods. We redefine the taxonomy of these domains, identifying the key domains of LLM agents for RSs as user interaction, representation optimization, system integration, and environment simulation.
Figure 3. Four domains of LLM agent’s role in recommendation tasks
Figure 3 shows the four domains used by LLM agents for recommendation tasks, with detailed descriptions of each domain as follows:
  • User Interaction: LLM agents serve as intermediaries between users and systems, enhancing interaction through natural language understanding and generation.
  • Representation Optimization: The innovation focuses on better understanding and representation of users/items through LLM agents, making the recommendation process more precise.
  • System Integration: LLM agent acts as the brain of RS, helping the RS analyze and make decisions for more effective results.
  • Environment Simulation: LLM agent is used as a user simulator to build a simulation test environment for the RS and optimize feedback to achieve continuous improvement.
We categorize research work according to the descriptions of the domains as mentioned above and list typical works related to the four domains in Table 1.
Table 1. A list of representative works of LLM agents for Recommendation.
Domain Paper What agents can do (ability)
Interaction RAH (Shu et al., 2024) Assists users in receiving customized recommendations and provide feedback
Interaction ToolRec (Zhao et al., 2024b) Uses tools for specific recommendation tasks
Interaction RecAI (Lian et al., 2024) Utilizes LLMs as an interface for traditional recommendation tools
Interaction AutoConcierge (Zeng et al., 2024) Conducts real conversations with users
Interaction FLOW (Cai et al., 2024) Introduces a feedback loop to enable collaboration between the recommendation agent and the user agent
Representation AgentCF (Zhang et al., 2024e) Facilitates collaborative learning between user and item agents
Representation Rec4Agentverse (Zhang et al., 2024a) Controls the collaboration between the Intelligent Agent items and the Agent Recommenders
Representation KGLA (Guo et al., 2024) Improves user agent memory
System RecMind (Wang et al., 2023a) Introduces a self-inspiring algorithm for decision-making
System InteRec (Huang et al., 2023a) Integrates LLMs and RSs for interactive recommendations
System MACRec (Wang et al., 2024d) Develops a multi-agent collaboration framework for RSs
System BiLLP (Shi et al., 2024) Emphasizes long-term user retention using LLM-planned RL algorithms
System MACRS (Fang et al., 2024) Tackles dialog control and user feedback integration with multi-agent framework
System PMS (Thakkar and Yadav, 2024) Uses multimodal, autonomous, multi-agent systems
System CORE (Jin et al., 2023) Combines conversational agents and RSs for better interaction
System Hybrid-MACRS (Nie et al., 2024) Combines LLM agent and search engine to optimize conversational recommendation
Simulation Agent4Rec (Zhang et al., 2023) Trains LLM agents to simulate real users for evaluation
Simulation RecAgent (Wang et al., 2023d) Simulates user behaviors related to the RS
Simulation Yoon et al. (Yoon et al., 2024) Uses LLMs to simulate users for conversational recommendation tasks
Simulation SUBER (Corecco et al., [n. d.]) Develops an RL environment using LLM to simulate user feedback
Simulation CSHI (Zhu et al., 2024b) Proposes a framework for LLM-based user simulators in conversational RSs
Simulation iEvaLM (Wang et al., 2023b) Suggests new evaluation methods using LLMs
Simulation Zhu et al. (Zhu et al., 2024a) Examines reliability and limitations of current LLM-based simulators
Simulation OS-1 (Xu et al., 2024b) Develops an LLM-based eyewear system with conversational common ground
Simulation CheatAgent (Ning et al., 2024) Uses LLM agent to attack LLM-driven RSs
Simulation Zhang et al. (Zhang et al., 2024f) Improves the training efficiency and effectiveness of RSs based on reinforcement learning
4.1.2. User Interaction
Traditional RSs rely on static data (such as user historical behavior) to model user preferences, but this approach is difficult to cope with scenarios where user preferences change rapidly, new users or new products (cold start problem), and niche product recommendations (long tail effect). In addition, although the dialogue recommendation algorithm in the traditional recommendation algorithm can dynamically capture user intentions through natural language interaction, its implementation faces challenges such as how to efficiently understand complex human language, how to maintain coherence and goal orientation in the dialogue, and how to deal with ambiguity in the dialogue. LLM agent supports users to interact in natural language and provides highly personalized recommendations by analyzing user preferences, historical behaviors, conversation context, etc. The LLM agent can also flexibly call external tools to further optimize recommendation results according to specific needs, providing users with a richer and more personalized interactive experience.
Specifically, AutoConcierge (Zeng et al., 2024) uses LLM to convert user questions into logical predicates, checks the consistency of information through the Answer Set Programming (ASP) system and updates the system status, and automatically generates questions to complete missing information when necessary. After obtaining all user preferences, the system searches for matching restaurants in its knowledge base and uses LLM to convert the results into natural language recommendations. If the user changes his mind or requests different suggestions, AutoConcierge can dynamically adjust its strategy, re-evaluate, and provide new recommendations. This process ensures efficient and accurate responses to user needs. ToolRec (Zhao et al., 2024b) framework uses LLM as a surrogate user to evaluate the degree of match between the user’s preferences and the current scenario, aiming to simulate the real user decision-making process. Then, external tools (ranking tools and retrieval tools) are called according to the user’s attribute instructions to explore different parts of the project pool. RecAI (Lian et al., 2024) is a practical toolkit that enhances or innovates RSs through the capabilities of LLMs. The LLM agent first develops a comprehensive execution plan based on the intent in the user conversation, then calls the tools, tracks the output of each tool, and ultimately generates a response for the user. In the FLOW (Cai et al., 2024) framework, the recommendation agent uses LLM, combined with the memory module and the recommendation module, to generate initial recommendations and optimize the recommendation results based on user feedback; the user agent is responsible for simulating user behavior and more accurately capturing the user’s potential interests by analyzing the interaction history with the recommendation agent. This iterative refinement process progressively enhances the cognitive capabilities of both the recommendation agent and the user simulation agent, consequently enabling more nuanced recommendation generation and higher-fidelity user behavior emulation. The RAH framework (Shu et al., 2024) combines RSs, assistants, and humans, using LLM agents to perceive, learn, act, criticize, and reflect. These agents collaborate through a learning-act-critic cycle to continuously improve their understanding of user personality. For example, when a user interacts or gives feedback, the learning agent extracts preliminary personality features, the action agent predicts user behavior based on this, and the criticism agent evaluates the accuracy of the prediction. If the prediction is inaccurate, the criticism agent will analyze the reasons in depth and make suggestions for improvement, and the learning agent will adjust the personality features accordingly until the prediction is consistent with the user’s actual behavior.
In summary, LLM agents can not only provide more accurate and personalized recommendation services but also provide users with a richer and more efficient interactive experience through continuous self-optimization and calling external resources. This marks the transformation of the RS from a simple information provider to an intelligent interactive partner, significantly improving the user experience and service quality.
4.1.3. Representation Optimization
In the traditional RS framework, the information records of each user and item are independent and static, including everything from basic information to detailed attributes, and most of these data updates rely on manual updates. However, user interests change dynamically, and different users have significantly different interests in the same item, and even the same user’s interests change at different times. Traditional representation methods based on static data have difficulty capturing these personalized needs and changes over time, which poses a challenge to providing accurate and effective recommendation services. The LLM agent can generate detailed representations through deep semantic analysis, multimodal data fusion, and external world knowledge, and it can keenly perceive context changes and user feedback, promoting dynamic learning and updating of user and item representations.
Specifically, AgentCF(Zhang et al., 2024e) treats both users and items as agents, each with its memory module to maintain the preferences and tastes of potential adopters. In addition, the study introduces a new collaborative learning method for simultaneously optimizing user agents and item agents and designs a collaborative reflection mechanism that enables agents to adjust their memories based on differences in real-world interaction records to more accurately reflect user behavior. KGLA (Guo et al., 2024) converts the path information in the knowledge graph into natural language descriptions, enhances the language agent’s understanding of user preferences, and dynamically updates user memory based on the interaction between users and items during the simulation phase. At the same time, it uses information such as brands, categories, and product features in the knowledge graph to generate detailed item descriptions, and more accurately characterizes item features by analyzing the path relationship between users and items, thereby improving the personalization and accuracy of the recommendation. In the Rec4Agentverse (Zhang et al., 2024a) framework, items are converted into interactive, intelligent, and proactive LLM agents that can dynamically acquire user preferences and continuously update their own characteristics through multiple rounds of dialogue. Unlike traditional static items, LLM agents can provide multi-dimensional knowledge representation and enhance service content through collaboration with other agents. For example, a travel agent can not only recommend itineraries based on user interests but also cooperate with other agents to provide more comprehensive services. In addition, agent items can accurately model user preferences and improve the accuracy of personalized recommendations and user experience.
These works effectively enrich the representation of users and items by leveraging individual agent autonomy and inter-agent collaboration, giving them more initiative. This shift makes recommendation algorithms more situational and dynamic and can capture subtle changes in user preferences and the potential value of product attributes. The domain is still in its infancy, with less relevant work, and more excellent results are expected to emerge in the future.
4.1.4. System Integration
Traditional RSs often use a batch processing architecture to regularly update models and recommendation results. For example, models can be retrained every day or every week to reflect the latest user preferences. Data cleaning, feature engineering, model selection, and so on need to be considered separately during system integration. In contrast, LLM agents significantly reduce the need for manual intervention by learning and dynamically adjusting recommendation strategies in real-time and supporting an instant feedback mechanism, which not only improves the efficiency of the recommendation system but also enhances the user experience. In addition, LLM agents also support a team collaboration framework where each agent can focus on different data sources or tasks. Through multi-level and multi-dimensional collaboration and information sharing between these agents, the system can more comprehensively understand user preferences and provide highly personalized and context-related recommendation content. From the perspective of framework structure, these studies can be divided into single-agent work (Wang et al., 2023a; Huang et al., 2023a; Jin et al., 2023; Shi et al., 2024), multi-agent collaboration (Fang et al., 2024; Wang et al., 2024d; Nie et al., 2024).
In the single-agent work category, RecMind (Wang et al., 2023a) introduces a self-inspiration algorithm for the LLM agent which is designed for RS to retain the status of all historical paths and use this historical information to optimize planning decisions, with the aim of addressing the limitations of existing RSs in generalizing their ability to perform new tasks and effectively utilizing external knowledge. InteRecAgent (Huang et al., 2023a) takes LLM as its core, processing instruction understanding, common sense reasoning, and human-computer interaction, while the recommendation model serves as a tool for domain knowledge and user behavior patterns. Through memory components and dynamic examples to enhance task planning and reflection mechanisms, it upgrades traditional RSs, such as ID-based matrix decomposition to systems that support natural language interaction. The CORE framework (Jin et al., 2023) adopts an offline training and online checking model, where the RS model acts as an offline relevance score estimator, while the LLM conversational agent checks these scores online to reduce uncertainty by minimizing the sum of unchecked relevance scores. The LLM agent in BiLLP (Shi et al., 2024) is designed as a high-level planner that divides the learning process into two levels: macro-learning and micro-learning. Macro learning is responsible for acquiring high-level guiding principles, while micro-learning focuses on learning personalized recommendation strategies. This dual-level approach aims to achieve long-term recommendation strategies in RSs.
In the multi-agent work category, MACRS (Fang et al., 2024) is a multi-agent conversational RS that controls dialogue flow, generates responses through multiple LLM agents, learns from user feedback to refine dialogue strategies, prevents errors, and interprets implicit user semantics. MACRec (Wang et al., 2024d) is a multi-agent framework that includes specialized agents like managers, analysts, reflectors, searchers, and task interpreters to address diverse recommendation tasks collaboratively. PAS (Thakkar and Yadav, 2024) consists of three agents: The first agent recommends products suitable for answering a given question; the second agent asks follow-up questions based on images belonging to these recommended products; and the third agent then conducts the autonomous search. It also features real-time data extraction, recommendations based on user preferences, and adaptive learning. Hybrid-MACRS (Nie et al., 2024) consists of a central agent and a search agent. The central agent is driven by LLM and is responsible for user interaction, preference identification, search query generation, recommendation generation, etc. The search agent consists of a search engine and a relative search module, which is responsible for processing the search queries generated by the Central Agent, returning a list of matching products, and optimizing search results using personalized sorting based on user historical behavior.
These investigations collectively strive to transcend the constraints inherent in conventional recommendation systems by harnessing the advanced capabilities of LLM agents across multiple dimensions, including natural language comprehension, personalized preference modeling, and context-aware adaptive decision-making.
4.1.5. Environment Simulation
In the past, collecting user feedback to optimize recommendation algorithms was a time-consuming and resource-intensive process. However, with the advancement of technology, a LLM agent has emerged as an advanced simulator that can simulate user behaviors and preferences and support real-time response and large-scale concurrent testing. By leveraging the highly simulated environment provided by LLM agents, development teams can now test and optimize recommendation algorithms more efficiently. This simulation not only greatly shortens the testing cycle, but also reduces the reliance on real user data, while ensuring user privacy and data security. Most of the studies (Wang et al., 2023b, d; Zhu et al., 2024b; Zhang et al., 2024b; Yoon et al., 2024; Zhu et al., 2024a) focus on using agents in conversational recommendation scenarios to simulate user behaviors and evaluate their effectiveness. In addition, some studies (Corecco et al., [n. d.]; Zhang et al., 2024f; Ning et al., 2024) have explored in depth the new challenges and risks that LLM agents may bring in more practical applications.
Specifically, iEvaLM (Wang et al., 2023b) represents an innovative interactive evaluation framework leveraging LLMs, incorporating an LLM-based user simulator to comprehensively model diverse system-user interaction scenarios. Extensive experimental evaluations conducted on two publicly available Conversational Recommender System (CRS) datasets demonstrate that iEvaLM achieves substantial performance enhancements over conventional evaluation methodologies. Notably, this framework introduces a novel dimension by incorporating the assessment of recommendation interpretability, thereby addressing a critical aspect often overlooked in traditional evaluation paradigms. Agent4Rec (Zhang et al., 2024b) generates 1,000 LLM-enabled agents with different social traits and preferences by initializing from the MovieLens-1M dataset. The agents interact with the RS on a page-by-page basis, performing operations such as watching, rating, evaluating, exiting, and interviewing. Experimental results show that the generative agent can effectively identify and respond to items that match user preferences, and the feedback provided by the agent can be used as enhanced data for iterative training and improvement of the recommendation strategy, thus forming a dual-track approach for comprehensive evaluation of recommendation algorithms. Yoon et al. (Yoon et al., 2024) introduced a novel evaluation framework designed to quantitatively assess the behavioral alignment between LLMs and human users in CRS. This comprehensive protocol encompasses five distinct evaluation tasks: (1) item selection for discussion, (2) binary preference expression, (3) open-ended preference articulation, (4) recommendation solicitation, and (5) feedback provision. Through systematic experimentation, their study not only identifies significant behavioral discrepancies between LLM-generated responses and human interactions but also provides empirical insights into mitigating these deviations through optimized model selection criteria and advanced prompting techniques. Zhu et al. (Zhu et al., 2024a) conducted an analysis of the inherent limitations in current LLM-based user simulation approaches, particularly focusing on the issue of inadvertent data leakage. To address these challenges, they proposed SimpleUserSim, an innovative framework that implements a controlled information disclosure mechanism. This novel simulator employs a direct conversational strategy to steer dialogue topics toward target items while maintaining strict information constraints - specifically, it restricts access to target item titles and operates solely on item attribute information until successful recommendation completion. The RecAgent (Wang et al., 2023d) framework has a detailed definition of LLM agent simulation users, including the Profile Module (ensuring that the agent has a unique personality), Memory Module (remember past behaviors and evolve dynamically in the environment), and Action Module (including RS-related searches, browsing, clicks, and page turns, as well as some social behaviors), so that it can effectively simulate real human behavior. The CSHI framework (Zhu et al., 2024b) introduces control mechanisms, enhances the scalability of simulation, and incorporates elements of human-computer collaboration, enabling agents to flexibly adjust their behavior based on real-time context and historical data, thereby providing a more accurate personalized experience in conversational RSs. Beyond acting as user simulators in conversational recommendation scenarios, LLM agents have further broadened their application areas, playing a vital role in reinforcement learning and attacks on RS. In the latest research, the SUBER framework (Corecco et al., [n. d.]) ingeniously utilizes LLM agents to simulate human behavior, thereby constructing a synthetic environment specifically for the training and evaluation of reinforcement learning-based RSs. This innovative approach effectively mitigates the high training costs associated with the scarcity of online data and addresses the challenges in evaluating RS models, particularly in accurately measuring model performance without directly impacting the real user experience. Zhang et al. (Zhang et al., 2024f) addresses the shortcomings of current user simulators such as SUBER and Agent4Rec (including high computational cost, susceptibility to “hallucinations”, and failure to fully capture the complex user-system interaction dynamics) by explicitly modeling user preferences, combining the advantages of logical reasoning and statistical learning, and designing an effective evaluation framework to improve the reliability and efficiency of the simulators. CheatAgent (Ning et al., 2024) first analyzes the parts of the text that have the most significant impact on recommendation results. Then, it utilizes an LLM agent to generate adversarial perturbations that can mislead the RS and optimize the attack strategy through prompt tuning. Finally, it fine-tunes the LLM’s attack strategy based on feedback from the target system to enhance the attack’s effectiveness.
In summary, these studies range from simple dialogue interactions to complex reinforcement learning environments, collectively representing the advancing application landscape of LLM agents in simulating user behavior.
4.2. LLM Agents for Search
4.2.1. Taxonomy Introduction
Typically, the interaction between users and search engines can be divided into four key steps: user intent, query writing, search execution, and result understanding (White, 2024). Figure 4 shows the five domains utilizing LLM agents in search tasks, which can correspond to the typical IR interaction process mentioned 2. Detailed descriptions of each domain are as follows:
Figure 4. Five domains of LLM agent’s role in search tasks
  • Task Decomposers: LLM agents can break down complex tasks into smaller, manageable components to align with user intent and enhance overall task execution. In this case, LLM agents serve as the leading interfaces of search engines.
  • Query Rewriters: LLM agents are good at refining user queries to make them clearer and more specific, leading to more accurate search results. In this case, LLM agents serve as the query modeling modules.
  • Action Executors: LLM agents can interact with various tools and APIs to gather necessary information on behalf of the users. In this case, LLM agents serve as the matching and ranking modules.
  • Results Synthesizers: LLM agents are suitable for summarizing large amounts of search results and helping users quickly grasp the essential details to make decisions. In this case, LLM agents bring new abilities into the traditional searching process.
  • User Simulators: LLMs can serve as user simulators to mimic user action and interaction for search systems improvement or evaluation. In this case, LLM agents give feedback to search engines.
We categorize research work related to search tasks according to the descriptions in the domains mentioned above and list typical works in the five domains in Table 2.
Table 2. A list of representative works of LLM agents in search.
Role of agent Paper What agents can do (ability)
Decomposer Laser (Ma et al., 2023b) Uses state-space exploration for web navigation tasks
Decomposer Knowagent (Zhu et al., 2024c) Integrates knowledge base for task decomposition and logical action execution
Decomposer Deng et al. (Deng et al., 2024c) Utilizes self-reflection memory enhancement planning for web navigation tasks
Decomposer Gur et al. (Gur et al., 2023a) Learns from experience to complete tasks and divide complex instructions
Decomposer SteP (Sodhi et al., 2023) Introduces dynamic strategy combination through task decomposition
Decomposer Koh et al. (Koh et al., 2024b) Enhances web navigation using tree search algorithms
Decomposer Agent Q (Putta et al., 2024) Integrates MCTS-guided search with self-critique for multi-step reasoning
Rewriter CoSearchAgent (Gong et al., 2024) Enables collaborative search through plug-ins that understand and refine queries
Rewriter Joko et al. (Joko et al., 2024) Assists in constructing personalized dialogue datasets to enhance query quality
Rewriter Aliannejadi et al. (Aliannejadi et al., 2024) Utilizes internal knowledge of LLMs for better retrieval and response generation
Rewriter Jain et al. (Jain et al., 2024) Proposes RAG-powered agents with multi-stream ensemble for semantic code search
Executor AVATAR (Wu et al., 2024a) Utilizes a comparator LLM to teach the agent how to use tools
Executor EASYTOOL (Yuan et al., 2024) extracts key information from tool documentation and designs a unified interface
Executor CodeAct (Wang et al., 2024a) Integrates LLm agents with a Python interpreter in order to execute code actions
Executor CodeNav (Gupta et al., 2024) Proposes code-as-tool paradigm through semantic code search engines
Synthesizer PersonaRAG (Zerhoudi and Granitzer, 2024) Uses real-time personalized data to enhance the relevance of the returned results.
Synthesizer ChatCite (Li et al., 2024a) Mimics human methods to extract key points and write summaries for literature reviews.
Synthesizer PaSa (He et al., 2025) Utilizes a selector to determine whether search results should be included or not.
Simulator Sekuli et al. (Sekulić et al., 2024a) Explores user emulators in conversation search systems for multi-round clarification
Simulator Usimagent (Zhang et al., 2024g) Simulates users’ query, click, and stop behavior in search tasks
Simulator BASES (Ren et al., 2024) Establishes a parameterized user profiling system with validation framework
Simulator Chatshop (Chen et al., 2024b) Introduces LLM-simulated shoppers to evaluate agents’ multi-turn interaction
4.2.2. Task Decomposers
The existing search technologies typically focus on learning the relevance between user queries and documents without any thoughtful processing. This limitation is critical for complex search tasks. For example, when a user searches for “where to go on vacation?”, his/her expectation might be a comparison of various travel destinations, details on travel itineraries, and hotel/flight prices. As discussed in Section 3, LLM agents possess strong reasoning and task decomposition abilities, enabling them to collect and aggregate information step by step and provide comprehensive results to the user.
Laser (Ma et al., 2023b) introduces a method that views web navigation tasks as state-space exploration. LLM agents decompose tasks into a predefined set of states to handle unfamiliar scenarios and flexible backtracking. Deng et al. (Deng et al., 2024c) propose a new task called Conversational Web Navigation, which requires complex multi-turn interactions with both users and the environment. In cases where tasks involve intricate actions, the agents will initially decompose the task into smaller sub-tasks. The decomposition results will later be verified and corrected by human annotators. Gur et al. (Gur et al., 2023a) introduce a LLM agent named WebAgent. WebAgent has the capability to decompose instructions into standardized sub-instructions and plan ahead, which enables it to better understand and execute complex tasks. For instance, in a complex web operation task, it can first identify the main steps that need to be executed and then carry out these steps sequentially. SteP (Sodhi et al., 2023) proposes a policy stack to better coordinate different actions and accomplish complex tasks. By decomposing tasks into various policy layers, each policy can focus on specific aspects of the task, thereby improving efficiency and accuracy in task execution. For example, in handling web tasks, one policy might be responsible for page navigation, while another policy might handle form filling. Together, these policies form a policy stack that collaboratively completes the entire task. Koh et al. (Koh et al., 2024b) propose an inference-time search algorithm where agents do a tree search to decide which action to take according to the current state. This method helps LLM agents better perform multi-step planning in interactive web environments. Knowagent (Zhu et al., 2024c) utilizes an action knowledge base and a knowledge-driven self-learning approach to guide the action paths during the planning phase. The action knowledge base furnishes the agent with extensive action knowledge, allowing for more rational action selection during planning. Meanwhile, the knowledge-driven self-learning approach enables the agent to continuously learn and refine its planning strategies, enhancing overall planning performance. Xie et al. (Xie and Zou, 2024) introduce a human-like reasoning framework to improve the planning ability that can help LLM agents when solving multi-phase problems like travel planning. Agent Q (Putta et al., 2024) proposes a framework that integrates guided Monte Carlo Tree Search (MCTS) with iterative fine-tuning for agents. This approach enables LLM agents to learn effectively from all trajectories, including both successful and unsuccessful ones, thereby enhancing their ability to generalize in complex, multi-step reasoning tasks.
Overall, the task decomposition capability of LLM agents significantly enhances the efficiency of users in obtaining information for complex search tasks. Some of the existing approaches predefine the action space and do search progress to decide the next move. These approaches offer strong control but are limited to solving specific problems. In contrast, the others allow the model to autonomously plan, offering greater flexibility, while the decomposition results may be in turn.
4.2.3. Query Rewriters
In traditional search scenarios, users must decide on the search terms themselves and master complex syntax to emphasize which keywords should necessarily be included or excluded from the search, and the length and context of these queries are usually limited. These limitations can prevent users from fully expressing their true intentions. However, these obstacles can be solved through LLM agents. LLM agents allow users to input longer context to fully express their intentions and then formulate more effective queries by themselves, thereby enhancing search efficiency.
CoSearchAgent (Gong et al., 2024) is a lightweight collaborative search agent powered by LLMs and serves as a search plugin on the instant message platform Slack. This agent can autonomously extract the user’s search intent based on the conversation between users, formulate it into a query, and then invoke a search engine to return the results. The entire process does not require the user to explicitly input a query, thereby avoiding disruption to the conversational experience. Joko et al. (Joko et al., 2024) propose a scheme called LAPS (LLM-Augmented Personalized Self-Dialogue) to gather personalized dialogues, which can be used to train preference extraction and personalized response generation. As more conversational data is collected, LLM agents can gradually learn about users’ various preferences. For instance, when a vegetarian user inquires about recipes, the agents can account for vegetarian needs and directly search for vegetarian-related rather than normal recipes. Aliannejadi et al. (Aliannejadi et al., 2024) organize a competition on how to utilize users’ prior interactions and present context and observed two main pipelines, namely, multi-turn retrieve-then-generate (R→G) and generate-retrieve-then-generate (G→R→G). In this context, “G” includes conversational query expansion and conversational rewriting. Jain et al. (Jain et al., 2024) propose an approach utilizing Retrieval Augmented Generation (RAG) to enable LLM agents to enrich user queries with additional information. Specifically, by leveraging RAG, the agents can augment user queries with pertinent details sourced from GitHub repositories.
In summary, LLM agents can significantly enhance users’ ability to utilize search systems. They can eliminate the need for users to learn search syntax, obviate the necessity of explicitly writing search queries, and even supplement information based on historical interactions or external data to improve search accuracy.
4.2.4. Action Executors
The searching process often involves the use of numerous tools. For example, in travel planning, users need to gather various types of information, such as geographical locations, hotels, flights, and weather. Specially designed LLM agents can learn how to judiciously select the appropriate tools and execute these actions on behalf of the user.
AVATAR (Wu et al., 2024a) is a novel automatic framework to optimize the performance of LLM agents when using external tools and knowledge to complete complex tasks. It consists of an actor LLM and a comparator LLM. During optimization, the actor generates actions to answer queries by utilizing the provided tools. The comparator then evaluates a set of well-performing (positive) and poorly-performing (negative) queries, teaching the actor more effective retrieval strategies and tool usage. EASYTOOL (Yuan et al., 2024) extracts key information from extensive tool documentation from various sources and meticulously designs a unified interface called tool instructions. This interface provides LLM-based agents with standardized tool descriptions and functionalities. By doing so, it effectively reduces the cognitive load on LLMs when understanding tool functionalities, thereby enhancing the efficiency and accuracy of tool usage. CodeAct (Wang et al., 2024a) proposes consolidating LLM agents’ actions into a unified action space using executable Python code with a Python interpreter integrated. In multi-turn interaction scenarios, CodeAct can execute code actions, dynamically revise previous actions, or generate new actions based on new observations. CodeNav (Gupta et al., 2024) points out that tool-using LLM agents typically require manual registration of all relevant tools in the LLM context, which presents limitations when dealing with complex real-world codebases. CodeNav addresses this issue by automatically indexing and searching code fragments within the target codebase, eliminating the need for manual tool registration. It can locate relevant code snippets and incorporate them to iteratively develop solutions. Compared to traditional tool-using LLM agents, CodeNav is more flexible and efficient, capable of adapting to different codebase structures and requirements.
In summary, recent researches have focused on two key areas. Firstly, how to equip LLMs agents with a broader range of tools, such as code execution and web retrieval. Secondly, how to make LLM agents better select the most appropriate tools from various options.
4.2.5. Results Synthesizers
In traditional search processes, after the search engine returns results, users must review each document individually, determine their relevance, and synthesize conclusions on their own. However, this laborious process can be entirely replaced by LLM agents, which can significantly enhance search efficiency and user experience.
PersonaRAG (Zerhoudi and Granitzer, 2024) is an innovative framework that incorporates user-centric agents to select retrieval and generation based on real user preference. This framework works through three steps: retrieval, user interaction analysis, and cognitive dynamic adaptation. PersonaRAG outperforms baseline models, delivering personalized answers that better address user needs on various question-answering datasets. ChatCite (Li et al., 2024a) is an LLM agent guided by human workflows for comparative literature summarization. The agent mimics the human workflow to extract key elements from relevant literature and generate summaries through a reflective incremental mechanism. This approach enables ChatCite to achieve a deeper understanding of the literature content and produce more targeted and comparative summaries. PaSa (He et al., 2025) introduces a hierarchical architecture comprising Crawler and Selector agents to address complex academic search challenges. The Crawler agent dynamically generates search queries and expands citation networks through tool invocation, while the Selector agent evaluates paper relevance via fine-grained scoring. To optimize this process, a novel conversational PPO algorithm is developed to handle sparse rewards in long-horizon search tasks.
In conclusion, large model agents can assist in summarizing search results from multiple perspectives. On one hand, they can consider personalized user information to filter and highlight content relevant to the user. On the other hand, they can be equipped with better mechanisms for filtering information, such as mimicking human expertise or employing specialized selectors.
4.2.6. User Simulators
Evaluating search effectiveness has always been a challenge. Existing solutions typically measure algorithm performance by conducting A/B tests and collecting the metrics difference between A/B groups, such as users’ click rate and stay time. However, these statistical measures cannot directly reflect user satisfaction, and conducting A/B tests may disrupt real user experiences when the test strategy is suboptimal. LLM agents offer a solution to these issues by simulating user behavior. As previously mentioned, with their independent reasoning and action capabilities, LLM agents can mimic user decision-making processes by learning from real user data. This allows them not only to provide simple feedback like clicks but also to articulate specific reasons for their preferences or dislikes.
Usimagent (Zhang et al., 2024g) is a search user behavior emulator based on LLMs. USimAgent can simulate users’ query, click, and stop behaviors in search tasks, generating complete search sessions. Through empirical research on actual user behavior datasets, researchers found that USimAgent outperforms existing methods in generating queries and performs similarly to traditional methods in predicting user click and stop behaviors. Sekuli (Sekulić et al., 2024a) explores the use cases of user simulators in conversation search systems. User simulators that can automatically answer clarification questions in multiple rounds of conversations and evaluate the rationality of system search results. Additionally, user simulators can generate large amounts of training data, thereby enhancing the performance of conversational systems. BASES (Ren et al., 2024) is an innovative user simulation framework based on LLM agents, designed to simulate web search user behavior comprehensively. It is capable of generating unique user profiles on a large scale, leading to diverse search behaviors. Specifically, through learning from vast amounts of web knowledge, the LLM can achieve human-like intelligence and generalization capabilities. Chen et al. (Chen et al., 2024b) proposed an online shopping task named ChatShop, where an agent needs to interact with the shopper to understand their needs and preferences gradually. They used LLM agents simulated shoppers to test the capabilities of the agent and found that the simulated shopper made the task more realistic, increasing its complexity and challenge like real human shoppers.
On the whole, the use of LLM agents as user simulators in conversational search systems shows great promise. They not only enhance evaluation efficiency but also offer valuable insights for improving these systems. Future research could explore further refinements and applications of such simulators to address the challenges in evaluating conversational search results.
4.3. Benchmark and Datasets
Existing benchmarks on LLM primarily focus on language understanding and generation. However, these datasets and benchmarks are inadequate for comprehensively assessing the agents on their decision-making and interaction capabilities in complex, dynamic environments. In this subsection, we will elaborate on the research approaches and progress in this area.
Furuta et al. (Furuta et al., 2024) introduced a benchmark, CompWoB, to highlight the limitations of language model agents in sequential task compositions and to emphasize the need for robust and generalized LLM agents for task combinations. CompWoB is a benchmark consisting of 50 new composite Web Automation tasks that reflect more realistic scenarios. Experimental results show that existing language model agents (such as gpt-3.5-turbo or gpt-4) achieve an average success rate of 94.0% on basic tasks, but their success rate drops to 24.9% on composite tasks. Agentbench (Liu et al., 2023) includes diverse agent-task scenarios and assesses LLM agent capabilities through task performance data. This benchmark not only pinpoints the strengths and weaknesses of LLMs in agent tasks but also provides a unified standard for comparison. WebArena (Zhou et al., 2023b) and MIND2WEB (Deng et al., 2024a) provide environments that replicate realistic web browsing experiences, incorporating fully functional websites from diverse domains, such as e-commerce and social forums, with URLs, open tabs, and keyboard and mouse interactions. Cocktail (Dai et al., 2024a) is a benchmark providing a comprehensive tool for evaluating information retrieval models in the LLM era. Cocktail comprises 16 different datasets covering various text retrieval tasks and domains, including a mix of human-written and LLM-generated corpora. The diversity of these datasets enables researchers to assess the performance of information retrieval models across different scenarios. To avoid potential biases from datasets previously included in LLM training, the authors introduce a new dataset named NQ-UTD. This dataset features queries originating from recent events, ensuring the fairness and validity of the evaluation.
Overall, researchers have introduced various datasets and benchmarks to evaluate the performance of LLM agents in IR scenarios. Unlike conventional LLM evaluation datasets that primarily focus on textual information, these datasets often include real online browsing tasks and user interaction behaviors.
5. Embodied LLM Agents: Towards Next Generation Recommendation and Search
Although LLM agents enhance traditional recommendation and search models with comprehensive capabilities, they are often limited to analyzing static user-item interactions, such as pre-training from fixed-size and limited-domain data. This limitation hinders their ability to respond to emerging interactions in changing environments such as smart devices, and provide proactive services from both user and environmental perspectives. Recent embodied agents take a forward step over the LLM agent above by actively perceiving and interacting with environments. Their applications in the cyber world, such as autonomous exploration and online content retrieval in web services, naturally encompass a myriad of information retrieval processes that closely align with the task of recommendation and search. In this section, we discuss the latest development of embodied agents in the cyber environment, their potential applications in recommendation and search, limitations, and promising directions.
5.1. Recent Advance of Embodied Agents in Cyber Environment
Embodied agents. LLMs have been recognized with notable reasoning abilities and capacities to utilize external tools and knowledge. Building on this foundation, researchers have integrated these capabilities into unified LLM agents with diverse domain profiles and knowledge of specific targets. However, these LLM agents remain constrained by their dependence on passive learning from static data of images, videos, and text, which are pre-sampled from real-world information systems, like the Internet.
In recent years, researchers have been exploring the next piece of the puzzle of realizing AGI. Reinforcement Learning (RL) and Embodied AI have emerged in response to the thriving research of LLM agents. Consequently, there has been a shift from “Internet agents” that focus on learning from constant datasets collected from the Internet, towards “Embodied agents” which enable LLM agents to learn through interactions with their surroundings (Duan et al., 2022). Just as its name implies, the term embodied grants the agent a physical substrate, manifesting in the real world as using robotic arms to measure and interact with the environment or in cyber environments as the “driving software” of information systems. Web agents serve as one of the key representations of embodied agents within the cyber environment. These agents must not only learn to navigate websites by interacting with GUI (Graphical User Interface) functions but also respond effectively to feedback from both web systems and human users. Such advancement inherently involves information retrieval tasks, requiring agents to perform complex multi-step searching and multi-modal retrievals. As a result, embodied agents are evolving into a new testbed and paradigm for research in recommendation and search areas.
How embodied agents interact with cyber environment. The current enthusiasm for building web agents is mainly divided into two directions; they consider static and dynamic environments respectively (Rawles et al., 2024a; Zhou et al., 2023b; Gur et al., 2023b). Most existing approaches follow the static environments by comparing an agent’s action trajectory to a pre-collected human demonstration. For example, AiTW (Rawles et al., 2024b), the most commonly used dataset, comprises over 700K sequential samples with screenshot-action pairs. This lightweight environment offers greater flexibility for designing model architectures and promotes the adoption of more complex technology stacks, such as Chain-of-Thought (CoT) (Deng et al., 2024b). However, agents trained within this constrained framework often follow a predetermined path to complete tasks, which can result in a limited understanding of real-world scenarios and difficulty in responding to new targets. To achieve more realistic evaluations, recent research has developed dynamic environments where agents learn interactively from their mistakes, enabling them to test the boundaries of these systems. AndroidWorld (Rawles et al., 2024a) is the representative benchmark where a simulated Android system is well-exhibited without restricted action trajectories. This introduces complexity and instability of the real environment, posing a huge challenge to the agent’s robustness and ability to cope with complex tasks. In summary, both of these directions entail extensive multi-modal retrieval tasks, which could represent either novel intersection fields or evaluation environments for recommendation and search research.
Applications. Along with these approaches, embodied agents have been developed in diverse fields related to recommendation and search areas. In website search, web agents have been used to instantiate search engine models into autonomous agents, equipped with external web-browsing tools (Koh et al., 2024a; Zheng et al., 2024a; He et al., 2024; Kim et al., 2024). In general GUI control, GUI agents have been supported by fine GUI element grounding capability to retrieve potential interactive objects (Hong et al., 2024; Zhan and Zhang, 2023; Ma et al., 2024; Cheng et al., 2024b). Other domains, such as games (Zhao et al., 2023a) and 3D environments (Huang et al., 2023b), have also witnessed the flourishing application of embodied agents.
5.2. Why Embodied Agents Can Be Used for Recommendation and Search
The widespread application of embodied agents is based on two key factors: the inherent human-like language reasoning ability of LLM and the interactive learning of embodied agents. The following discussion delves into the intrinsic qualities behind these factors that make embodied agents viable tools for recommendation and search tasks, by reviewing their capabilities as lifelong learners, one-for-all (One4all) models, and personal assistants, respectively.
  • Lifelong learner: One of the core strengths of embodied agents is their ability for continuous and even lifelong learning (Wang et al., 2023c). Unlike traditional machine learning models that are typically trained on a fixed dataset, embodied agents can persistently expand their knowledge and skills through ongoing interaction with environments. In the recommendation field, previous lifelong works (Yuan et al., 2021; Lin et al., 2023; Li et al., 2023d) continually focus on identifying long-term behavior dependencies and incorporating emerging intentions, while these traditional task-specific models constantly suffer from catastrophic forgetting. The nature of embodied agents as lifelong learners could be a novel solution.
  • One4all model: Another crucial advantage of embodied agents is their ability to achieve the One4all model, i.e., generalist problem solver. Differing from learning lifelong, the capability of one4all modeling concentrates on cross-domain tasks, where one4all agents are expected to efficiently adapt to unseen domains at minimal cost (in terms of adaptation time, fine-tuning data, loss of performance, etc.). Existing recommendation work (Fu et al., 2024a, b; Li et al., 2023b; Cheng et al., 2024a) has made great efforts on this topic thanks to techniques such as large-scale pre-training and efficient adaptation. However, a universal representation that seamlessly spans various domains is still out of reach (Shin et al., 2021; Li et al., 2023a; Zhang et al., 2024c). Embodied agents could contribute massive supportive information to knowledge transfer through rapid adaptation to new scenarios.
  • Personal assistant: Embodied agents can be omnipotent personal assistants. The characteristics of learning lifelong and being flexible as the one4all model could make the embodied agent a competent assistant for perceiving the rapidly changing personalized needs in the spatial-temporal dimension. In recommendation and search tasks, the possibility of condensing well-trained models into cloud-edge personal services gradually becomes a hot spot in both academics and industries. Prior research (Wang et al., 2024b; Yuan et al., 2023; Yin et al., 2024; Rawassizadeh and Rong, 2023) have delved into the path of achieving lightweight on-device recommender systems and search tools in privacy-preserving and cybersecure approaches, where embodied agents could be the next generation of solutions.
Embodied Agent Designing for Recommendation and Search. Existing embodied agents within cyber environment primarily focused on interaction with GUI. These agents concentrate on GUI operations and task-solving on a wide range of information systems like websites, desktops, and smartphones, which are typical scenarios for recommendation and search tasks. This close relationship could inspire innovative approaches and solutions for the next generation of recommendation and search. WorkArena (Drouin et al., 2024) introduces the first environment that supports chat-based agent-user interactions for the development and evaluation of web agents. Through a chat interface, the real human user can exchange messages with web agents. This allows for information retrieval tasks where a specific answer is expected from the agent, but also more practical tasks where user instructions change over time. This advancement may inspire further research and discussion around recommendation and search tasks, as agents would need to complete information retrieval tasks and output specific actions or plans. SmartAgent (Zhang et al., 2024d) defines a new task known as embodied personalized learning, where embodied cognition and item recommendation are interlaced as in many real-world personal assistance scenarios. To address this, a novel reasoning paradigm called Chain-of-User-Thought (COUT) is proposed to align user feedback with embodied agent actions under the progressive thoughts from basic GUI navigation to explicit and implicit user requirements. Leveraging the COUT paradigm, SmartAgent demonstrates the first full-stage embodied personalized capabilities in collaboration with a new dataset called SmartSpot, which for the first time supports a range of diverse embodied personalized environments.
In summary, the combination of embodied agents with recommender and search systems could lead to the creation of intelligent systems offering more personalized services, but may also realize truly bionic personal assistants. This could be a potentially important direction for the next generation of recommendation and search research.
5.3. Limitation and Promising Directions
Limitation. Embodied agents have shown promise in the fields of recommendation and search, but they are also challenged by several key limitations. One major challenge is the limited generalization capability of current agents. Most existing agents are primarily designed for static environments, resulting in suboptimal performance when faced with dynamic real-world scenarios. Additionally, these static environments typically represent one-sided functions of specific systems, which further complicates the training of general embodied agents across diverse situations. This further leads to the second issue, where currently embodied agents are still not ready to effectively handle information retrieval tasks involved in complex systems. This is mainly due to their limited understanding of environments, as they are still stuck on learning systems’ functionality. However, real-world application scenarios frequently involve personalized needs, which are often expressed through non-standard paths that extend beyond the training samples. Goal-completion-oriented embodied agents may struggle to address these personalized requirements underlying, resulting in lower user engagement and stickiness. Finally, there are still other critical technical bottlenecks that have not been fully addressed. For example, executing advanced cross-app operations on smartphone devices and multi-web page tasks on desktop environments remains challenging.
Promising directions. These limitations also spur new research directions. One of them is leveraging the one-for-all capability of embodied agents for transferable recommendation or search. Currently, large-scale pre-trained recommendation models require extremely high costs to transfer to new scenarios. In some new scenarios where the gap is relatively large, the current pre-training and fine-tuning (PEFT) paradigm finds it difficult to achieve better domain adaptation. The ability of embodied agents to understand dynamic changes in the environment may be able to help with this. Another promising direction is the use of lightweight embodied agents for on-device recommendation. LLMs are often too resource-intensive to deploy on mobile devices, hindering the implementation of personalized services. Embodied agents, with their smaller footprint, may provide a solution to this problem. Additionally, the demand for zero-shot services on end-user devices also poses new requirements for embodied agent capabilities, which could drive more research. Finally, as embodied agents extend the capabilities of LLMs to recommendation and search areas within personal settings, the issue of privacy preservation becomes more crucial. Balancing user behavior modeling and individual privacy protection while providing intelligent services will be a critical and urgent problem.
6. Open Problems and Future Directions
Even though LLM agents have shown great potential to bring promising advancement to IR, there are still many challenges to address and numerous directions to explore. In this section, we will discuss these topics.
Hallucinations. Hallucinations in LLMs refer to situations where the content generated by LLMs is inconsistent with facts or cannot be verified. For LLM agents in the IR field, hallucinations can result in the retrieval of incorrect information, thereby degrading the user experience. Hallucinations can be divided into factuality hallucinations and faithfulness hallucinations (Huang et al., 2023c). The former refers to the inconsistency between the content generated by the LLM and the verifiable facts of the real world, while the latter refers to the inconsistency between the content generated by the model and the user’s instructions. Researchers have proposed many methods to alleviate these two types of hallucinations. For factual hallucinations, solutions such as RAG (Gao et al., 2023a; Xu et al., 2024a; Ram et al., 2023), CoT (Wei et al., 2022), and consistency enhancing (Wang et al., 2022; Shi et al., 2023; O’Brien and Lewis, 2023) can constrain LLM to obtain data from given sources. For command-type hallucinations, prompt engineering, RLHF, and other alignment technologies (Sharma et al., 2023; Perez et al., 2022; Wei et al., 2023) can strengthen the consistency of execution.
Bias. Bias is a classic issue in the IR field, typically arising from improper data collection or training methods, leading to a significant discrepancy between the model results and real data, thereby degrading the accuracy of information retrieval. The capabilities of LLMs are largely constrained by their training data, making them more likely to encounter bias problems. Studies (Dai et al., 2023, 2024b; Ma et al., 2023a; Zheng et al., 2023) indicate that LLMs can introduce various biases, such as gender discrimination (Dhingra et al., 2023), political extremism (Dong et al., 2024), and echo chamber (Sharma et al., 2024). Dai et al. (Dai et al., 2023) find that LLM-based retrievers tend to prioritize content generated by LLMs over human-written text and introduce a bias-oriented loss to alleviate this trend.
Deployment Cost. LLM agents are typically composed of multiple LLM modules, resulting in a higher number of invocations and model parameters compared to general LLM applications. So, deploying LLM agents requires substantial resources, resulting in longer inference latency, which imposes higher hardware requirements on the end devices and may impact users’ experience. Researchers have put a lot of effort into this issue. Wilkins et al. (Wilkins et al., 2024) optimize LLM efficiency by proposing a cost-based scheduling framework, while Avatar (Wu et al., 2025) and QueryAgent (Huang et al., 2024) utilize tools or environmental feedback to reduce inference time.
Multi-Modal Agent. Multi-modal LLM agents can integrate data from different modalities, such as images and text, to more comprehensively understand user queries. For instance, when querying about a product, a user might provide both a textual description and an image of the product. A multi-modal LLM agent can analyze both the text and image information simultaneously, better understand the product’s features, and thereby provide more accurate retrieval results (Xie et al., 2024; Zhang et al., 2024h). Recently, numerous studies have begun exploring how to leverage multi-modal capabilities to enhance LLM agents’ ability to operate on web pages (Zheng et al., 2024b; Furuta et al., 2023) and mobile devices (Nong et al., 2024; Wang et al., 2024c). These studies typically utilize visual perception capabilities to identify visual and textual elements of mobile applications accurately.
Domain Specific Agent. In specialized domains with substantial domain knowledge, general LLM agents often perform poorly due to their reliance on common knowledge and data sources. However, data in these specialized domains is relatively sparse, making it insufficient to fully train LLMs. Therefore, efficiently and economically transforming general LLM agents into domain-specific LLM agents is a crucial direction for future research. Currently, related research has already been applied to relatively niche but popular categories such as clinical (Van Veen et al., 2024; Li et al., 2023c; Russe et al., 2024; Yan et al., 2024; Jin et al., 2024b) and legal (Sun et al., 2024; Li et al., 2024b). In the future, as LLM agents progressively improve their ability to utilize sparse data, there may be opportunities for enhancements in niche areas such as psychological counseling and job seeking.
Multi Agent Interaction. In multi-agent systems, individual agents can solve complex problems through information sharing and collaboration. Different agents may possess distinct knowledge and skills, and by cooperating, they can leverage their strengths to enhance overall performance information retrieval (Choi and Yoo, 1998; Wondergem et al., 1998). Recently, studies (Park et al., 2023, 2024) have explored the interaction and evolution processes among multiple LLM agents, and several multi-agent frameworks (Chen et al., 2023; Qian et al., 2023; Team, 2023) have been proposed. Infogent (Reddy et al., 2024) utilized multi-agent systems to simulate human cognition for gathering information from the web. Chateval (Chan et al., 2023) and MAD (Liang et al., 2023) improve agents’ generating quality through multi-agent debate.
Personalization. Different users often have varying needs and preferences when performing information retrieval. For instance, the search term “Apple” might refer to Apple Inc. for a tech investor, whereas it might mean the fruit for others. By leveraging the memory module and long context processing capabilities of LLM agents, personalized information retrieval tailored to individual users can be achieved (Li et al., 2024c). One approach to do so is to treat the user’s interaction history as context and input it into the model (Mo et al., 2024; Richardson et al., 2023; Aknouche et al., 2012). Mo et al. (Mo et al., 2024) incorporate conversational and personalized elements with LLM agents to fulfill user needs through multi-turn interactions. They indicate that Personal Textual Knowledge Bases (PTKBs) can effectively enhance conversational information retrieval, allowing retrieval results to align more closely with the user’s background.
7. Conclusions
In this paper, we take the pioneering step in reviewing the advanced applications of LLM agents in recommender and search systems. We begin the survey with a brief introduction to the concepts of recommendation and search tasks, as well as LLM agents, offering newcomers a foundational overview and critical background knowledge. Then, we elucidate the intrinsic nature of LLM agents, highlighting three specific aspects to underscore why LLM agents can be used for recommendation and search. Moreover, we provide a detailed introduction to how to apply LLM agents in each task. For each phase, we provide a detailed taxonomy to categorize the major techniques and roles, drawing connections among the existing publications. Additionally, we extensively discuss the potential of applying embodied agents in both tasks, delving into the rationale behind the advanced development. Finally, we summarize several challenges and promising directions in this field, which are expected to guide potential future directions.